# Adaptive finite element interpolated neural networks

### v1

@@ 1. Introduction
pde are an effective mathematical tool for modelling a variety of physical phenomena in science and engineering, such as heat conduction, fluid dynamics and electromagnetism. Since analytical solutions of partial differential equations are only available in few special cases, numerical methods, such as the finite element method (FEM), become the primary approach for approximating PDE solutions using computers. The FEM has solid mathematical foundations [1] and it can readily handle a broad range of PDE problems. Over the years, advanced discretisation techniques have been proposed [2], and highly efficient (non)linear solvers that can exploit large-scale supercomputers have been designed and polished [3, 4, 5].
The FEM is a linear method that provides the best approximation in some specific measure on a finite-dimensional vector space. However, this space is not adapted to local features (e.g., sharp gradients or singularities) and the convergence is slow for problems that exhibit multiple scales. Adaptive FEs, e.g., the $h$-adaptive FEM, adapt the space to the solution by gradually increasing resolution in regions with larger a posteriori error estimations. It adds an external loop, known as the solve, estimate, mark and refine loop, to the standard FEM workflow. Robust a posteriori error estimators have been designed for diverse problems (see e.g., [6, 7]), making the adaptive FEM a widely employed technique in scientific disciplines.
@@
Many recent works have tried to address the limitations of PINNs, aiming to improve their robustness, efficiency and accuracy. The authors in [14] introduce VPINNs, which make use of a polynomial test space and take the norm of the corresponding residual vector as the loss function. To overcome the integration challenge in VPINNs, the authors in [15] propose the interpolated (IVPINN) method. This method interpolates a NN onto a FE space and uses this polynomial interpolation in the variational residual. In PINNs, the imposition of Dirichlet boundary conditions is a challenging task; as opposed to FE spaces, NNs are not interpolatory objects. The addition of a penalty term to weakly impose the Dirichlet boundary conditions has a very negative effect on the training. In [15], the authors compose the NN with a polynomial function to design a NN that vanishes on the boundary. This approach is applicable to simple geometries but its extension to general geometries is unclear [16]. Besides, while better than the penalty method, such composition negatively impacts the generalisation of NNs themselves [17].
@@
Our previous work [17] introduces the interpolated (FEINN) method. This method interpolates the NN onto a FE space that strongly satisfies the boundary condition, avoiding the issues discussed above. Besides, we propose to use the dual norm of the PDE residual as the loss function, which can be re-stated as a preconditioned loss function that is well-posed at the continuous level; see also [18] for the application of dual norms in VPINNs. The preconditioning requires the application of the inverse of the inner product matrix (e.g., a Laplacian in $H^{1}(\Omega)$-conforming cases), which we efficiently approximate using geometric multigrid techniques. The FEINN method has been successfully applied to, e.g., forward PDEs with smooth or non-smooth solutions and inverse heat conduction design problems. We show that the trained NNs can outperform the FEM solution by several orders of magnitude when the exact solution is smooth.
@@
Despite these efforts, the training of PINNs and related methods often fail when dealing with target solutions containing high-frequency modes [19], e.g., solutions with sharp gradients or discontinuities. Some works aim to improve the expressivity of the NN to capture these solutions, adding more neurons/layers [20], or adapting the non-linear activation functions [21]. However, none of these works solve the integration issue; a quadrature on a uniform mesh is used in [20] and random collocation points in [21]. To solve this issue, several approaches propose an adaptive distribution of points in collocation-based PINNs. The authors in [22] design a residual-based adaptive refinement process to add more collocation points in the locations with the largest PDE residual, while a deep generative model that guides the sampling of collocation points is proposed in [23]. The adaptive collocation point movement method in [24] resamples the collocation points and uses an adaptive loss weighting strategy. In [25], in addition to the strong residual of the PDE, the gradient information of the NN is also used to guide the sampling procedure. All these adaptive methods have demonstrated enhanced performance and robustness compared to standard PINNs when solving PDEs with steep gradients. Still, they suffer the applicability and robustness issues of collocated PINNs mentioned above. The provably accurate adaptive quadratures in [26] address the integration issue of NNs by automatically creating an accurate quadrature rule for loss functions that involve the integration of PDE residuals. However, the computation of the quadrature is computationally expensive beyond 2D.
@@
A posteriori error estimations are crucial for adaptive methods. Since the NN in our method is interpolated onto a FE space, we can readily leverage the well-studied FE error estimators [6, 27]. In recent years, NN-based a posteriori error estimators have been developed. In [28], the authors propose a NN error estimator based on the dual weighted residual method for PINNs. Another related work is [29], where an error estimator for VPINNs consisting of a residual-type term, a loss-function term, and data oscillation terms is introduced. Functional error control techniques have been derived in [30] to estimate the error in deep Galerkin simulations.
@@
Remarkably, the state-of-the-art contains few investigations into the application of NNs for problems with singular solutions. In [15, 17], the authors use the FE interpolation of the trained NNs as the final solution of a 2D PDE with singularity. The deep Fourier residual (DFR) method [31] tackles problems with singular solutions using NNs via adaptive domain decomposition. However, the method is still prototypical in the sense that automatic refinement is currently limited to 1D scenarios and relies on tensor products of 1D orthonormal basis functions and an overkill integration procedure is used.
@@
In our previous study [17], we have demonstrated FEINNs outstanding performance in solving various forward and inverse problems. In this work, we carry out a detailed numerical analysis of the method for the FE residual minimisation in a dual norm. Since the training of FEINNs on a fixed underlying mesh limits their ability to leverage the full potential of NNs non-linear capabilities, we propose the $h$-adaptive FEINN method to overcome this limitation by introducing an automatic adaptation of the interpolation space (mesh) during training, driven by a posteriori error estimation. We propose a NN error estimator based on the strong PDE residual (which is straightforward to implement and can be accurately computed via Gaussian quadratures) and a standard FE Kelly indicator. These advances allow for a more comprehensive exploitation of the non-linear nature of NNs, thus enhancing the capability of the method to tackle problems with localised features. The minimisation of the loss function relies on a train, estimate, mark, adapt procedure, which is repeated till a certain level of accuracy is reached. We leverage hierarchically adapted non-conforming octree-based meshes and adaptive FE spaces with hanging node constraints to keep the conformity of the interpolation.
@@
We conduct a comprehensive set of numerical experiments, focusing on forward problems featuring steep gradients in 2D and singularities in both 2D and 3D. We verify that expressive enough NNs can replicate the rate of convergence of the $h$-adaptive FEM solutions. In addressing the sharp-gradient problem, the generalisation capabilities of the trained NNs yield remarkable results. The non-interpolated NN solutions can be orders of magnitude more accurate than their interpolation counterparts on the same mesh. This observation is consistent with the findings in our previous work [17] for smooth solutions. Our findings reveal that the performance of the NN error indicator is on par with, if not surpassing, the traditional Kelly FE error estimator in various situations. Moreover, we confirm that a well-posed residual norm in the loss function has a tremendously positive impact on the convergence of the $h$-adaptive training process.
@@ 2.2. Finite element method
Problem (5) is well-posed under the assumption that the trial and test FE spaces $U_{h}$, $V_{h}$ have the same dimension and satisfy the following inf-sup condition [32]: there exists a constant $\beta>0$ independent of the mesh size $h$ such that
@@ 2.4. Finite element interpolated neural networks
We refer to the solution of (11) as an $\epsilon$-quasiminimiser of the FEINN problem. The concept of quasiminimisers in the frame of PINNs has already been previously discussed [33, 34]. This is of relative practical relevance since the problem is solved iteratively up to a certain level of accuracy and one can produce a sequence of $\epsilon$-quasiminimisers that converges to one of the solutions of the FEINN problem (10).
@@
It has been proven in [35] that the space of functions that can be generated by a NN is not closed in $L^{p}$-norms for $0<p<\infty$ for most activation functions. For the $L^{\infty}$-norm, and (parameterised) ReLU activation functions, the NN space is proven to be closed in some cases. ReLU activation functions are not smooth enough for PINNs or Deep Galerkin methods [26]. However, in FEINNs, $\mathcal{N}_{h}$ only depends on the pointwise values of the neural network on the nodes of $U_{h}$ and ReLU activation functions are suitable, since the method computes spatial derivatives of FE functions. Thus, using the results in [35, Sec. 3.2], one can check that $\mathcal{N}_{h}$ is closed for ReLU activation functions and the minimum in (10) is attained. This is not the case for other activation functions (see Ex. 3.11).
As discussed in [17], we can define a discrete Riesz projector $\mathcal{B}_{h}^{-1}:V_{h}^{\prime}\to V_{h}$ such that
@@
Unlike loss functions that rely on discrete norms of the residual vector, this loss function, preconditioned by $\mathcal{B}_{h}$, is well-defined in the limit $h\downarrow 0$. In practice, $\mathcal{B}_{h}^{-1}$ can be replaced by any spectrally equivalent approximation to reduce computational costs; e.g. one geometric multigrid (GMG) cycle has been proposed in [17].
@@
where $\left\lVert\cdot\right\rVert_{\chi}$ is an algebraic norm of a vector. The algebraic norm is ill-posed at the continuous level and, as a consequence, the convergence of the minimisation process is expected to deteriorate as $h\downarrow 0$ [36].
@@ 2.6. Gradient-conforming discretisation
We consider an exotic Petrov-Galerkin FE method. The well-posedness of the discretisation is guaranteed by the discrete inf-sup condition (6) and the same dimension for trial and test spaces (see [32]). We use a grad-conforming nodal Lagrangian FE space $U_{h}$ of order $k_{U}$ built upon a (possibly non-conforming) partition $\mathcal{T}_{h}$ of $\Omega$. The FE space $U_{h}$ is built upon the partition $\mathcal{T}_{h}$. We use a linearised test FE space $V_{h}$ ($k_{V}=1$), built upon the partition $\mathcal{T}_{h}$ obtained after $k_{U}$ levels of uniform refinement applied to $\mathcal{T}_{h}$. It is easy to check that the dimension of ${U}_{h}$ and $V_{h}$ are identical. The reason behind this choice of the test space is the dramatic improvement in the convergence of the optimisation process compared to a Galerkin discretisation, as observed, e.g., in [15, 17].
@@ 2.7. Error indicators
The first error indicator is the classic Kelly error estimator [27, 6], a well-established indicator in $h$-adaptive FEM, particularly for the Poisson equation. This error indicator estimates the error per element by integrating the gradient jump of the NN interpolation along the faces of each element. Specifically, for the element $K$, the Kelly error indicator is defined as
@@
The superscript $n$ in $\zeta_{K}^{n}$ stands for “network”. We note that $\zeta_{K}^{n}$ is similar to the mean PDE residual $\varepsilon_{r}$ defined in [22]. The main difference is that $\zeta_{K}^{n}$ is evaluated with a high-degree Gauss quadrature, while $\varepsilon_{r}$ is computed by Monte Carlo integration. At each mesh adaptation step, the strategy in [22] adds more collocation points only on the region with the highest $\varepsilon_{r}$. Instead, our approach can refine multiple elements with the largest errors and coarsen those with the smallest errors simultaneously.
@@ 3. Numerical analysis
In this section, we extend the analysis in [17] in multiple ways. In particular, in Sec. 3.1, we bound the error of the interpolated NN under the general scenario in which the NN cannot exactly emulate the FE space $U_{h}$, i.e., $\mathcal{N}_{h}\subsetneq U_{h}$ (or, equivalently, the interpolation operator $\pi_{h}:\mathcal{N}\mapsto U_{h}$ is not surjective). Remarkably, we prove that the error of the interpolated NN is bounded by the interpolation error among all possible vector spaces that result from the interpolation of the last-layer functions in the NN architecture. In Sec. 3.2, we perform error analysis under a weakened assumption of exact emulation which we refer to as quasi-emulation. This analysis is relevant as exact emulation might not be feasible for some activation functions, regardless of the expressivity of the NN. Finally, in Sec. 3.3, we also discuss the well-posedness of the FEINN problem in a quotient space and discuss the accuracy of the non-interpolated NN by considering the maximum distance between elements of an equivalence class.
@@ 3.1. Error analysis
First, we note that the loss function is differentiable (see [17, Prop. 3.1]). To obtain (17)-(18), we compute the variational derivative of (16) to variations of $u_{h}$ and recall the definition of the discrete Riesz projector. Under the assumption of the discrete inf-sup condition (6), we can prove the well-posedness of the problem using the Babuska-Brezzi theory. Namely, using (6), we can readily find $k_{h}\in V_{h}$ such that $\|k_{h}\|_{U}=\|u_{h}\|_{U}$ and $a(u_{h},k_{h})\geq\beta\|u_{h}\|_{U}^{2}$. We can take $v_{h}=r_{h}+\alpha k_{h}$ for an arbitrary $\alpha>0$ and $w_{h}=-u_{h}$, to get

### v2

@@ 1. Introduction
pde are an effective mathematical tool for modelling a variety of physical phenomena in science and engineering, such as heat conduction, fluid dynamics and electromagnetism. Since analytical solutions of partial differential equations are only available in few special cases, numerical methods, such as the finite element method (FEM), become the primary approach for approximating PDE solutions using computers. The FEM has solid mathematical foundations [Ern2021] and it can readily handle a broad range of PDE problems. Over the years, advanced discretisation techniques have been proposed [Arnold2006], and highly efficient (non)linear solvers that can exploit large-scale supercomputers have been designed and polished [Brune2015, Badia2016, Drzisga2017].
The FEM is a linear method that provides the best approximation in some specific measure on a finite-dimensional vector space. However, this space is not adapted to local features (e.g., sharp gradients or singularities) and the convergence is slow for problems that exhibit multiple scales. Adaptive FEs, e.g., the $h$-adaptive FEM, adapt the space to the solution by gradually increasing resolution in regions with larger a posteriori error estimations. It adds an external loop, known as the solve, estimate, mark and refine loop, to the standard FEM workflow. Robust a posteriori error estimators have been designed for diverse problems (see e.g., [Ainsworth1997, Schoeberl2008]), making the adaptive FEM a widely employed technique in scientific disciplines.
@@
Many recent works have tried to address the limitations of PINNs, aiming to improve their robustness, efficiency and accuracy. The authors in [Kharazmi2021] introduce VPINNs, which make use of a polynomial test space and take the norm of the corresponding residual vector as the loss function. To overcome the integration challenge in VPINNs, the authors in [Berrone2022] propose the interpolated (IVPINN) method. This method interpolates a NN onto a FE space and uses this polynomial interpolation in the variational residual. In PINNs, the imposition of Dirichlet boundary conditions is a challenging task; as opposed to FE spaces, NNs are not interpolatory objects. The addition of a penalty term to weakly impose the Dirichlet boundary conditions has a very negative effect on the training. In [Berrone2022], the authors compose the NN with a polynomial function to design a NN that vanishes on the boundary. This approach is applicable to simple geometries but its extension to general geometries is unclear [Sukumar2022]. Besides, while better than the penalty method, such composition negatively impacts the generalisation of NNs themselves [Badia2024].
@@
Our previous work [Badia2024] introduces the interpolated (FEINN) method. This method interpolates the NN onto a FE space that strongly satisfies the boundary condition, avoiding the issues discussed above. Besides, we propose to use the dual norm of the PDE residual as the loss function, which can be re-stated as a preconditioned loss function that is well-posed at the continuous level; see also [Rojas2023] for the application of dual norms in VPINNs. The preconditioning requires the application of the inverse of the inner product matrix (e.g., a Laplacian in $H^{1}(\Omega)$-conforming cases), which we efficiently approximate using geometric multigrid techniques. The FEINN method has been successfully applied to, e.g., forward PDEs with smooth or non-smooth solutions and inverse heat conduction design problems. We show that the trained NNs can outperform the FEM solution by several orders of magnitude when the exact solution is smooth.
@@
Despite these efforts, the training of PINNs and related methods often fail when dealing with target solutions containing high-frequency modes [Wang2022], e.g., solutions with sharp gradients or discontinuities. Some works aim to improve the expressivity of the NN to capture these solutions, adding more neurons/layers [Cai2022], or adapting the non-linear activation functions [Jagtap2020]. However, none of these works solve the integration issue; a quadrature on a uniform mesh is used in [Cai2022] and random collocation points in [Jagtap2020]. To solve this issue, several approaches propose an adaptive distribution of points in collocation-based PINNs. The authors in [Lu2021] design a residual-based adaptive refinement process to add more collocation points in the locations with the largest PDE residual, while a deep generative model that guides the sampling of collocation points is proposed in [Tang2023]. The adaptive collocation point movement method in [Hou2023] resamples the collocation points and uses an adaptive loss weighting strategy. In [Mao2023], in addition to the strong residual of the PDE, the gradient information of the NN is also used to guide the sampling procedure. All these adaptive methods have demonstrated enhanced performance and robustness compared to standard PINNs when solving PDEs with steep gradients. Still, they suffer the applicability and robustness issues of collocated PINNs mentioned above. The provably accurate adaptive quadratures in [Magueresse2024] address the integration issue of NNs by automatically creating an accurate quadrature rule for loss functions that involve the integration of PDE residuals. However, the computation of the quadrature is computationally expensive beyond 2D.
@@
A posteriori error estimations are crucial for adaptive methods. Since the NN in our method is interpolated onto a FE space, we can readily leverage the well-studied FE error estimators [Ainsworth1997, Kelly1983]. In recent years, NN-based a posteriori error estimators have been developed. In [Minakowski2023], the authors propose a NN error estimator based on the dual weighted residual method for PINNs. Another related work is [Berrone2022_err_analysis], where an error estimator for VPINNs consisting of a residual-type term, a loss-function term, and data oscillation terms is introduced. Functional error control techniques have been derived in [Muzalevskiy2023] to estimate the error in deep Galerkin simulations.
@@
Remarkably, the state-of-the-art contains few investigations into the application of NNs for problems with singular solutions. In [Berrone2022, Badia2024], the authors use the FE interpolation of the trained NNs as the final solution of a 2D PDE with singularity. The deep Fourier residual (DFR) method [Taylor2024] tackles problems with singular solutions using NNs via adaptive domain decomposition. However, the method is still prototypical in the sense that automatic refinement is currently limited to 1D scenarios and relies on tensor products of 1D orthonormal basis functions and an overkill integration procedure is used.
@@
In our previous study [Badia2024], we have demonstrated FEINNs outstanding performance in solving various forward and inverse problems. In this work, we carry out a detailed numerical analysis of the method for the FE residual minimisation in a dual norm. Since the training of FEINNs on a fixed underlying mesh limits their ability to leverage the full potential of NNs non-linear capabilities, we propose the $h$-adaptive FEINN method to overcome this limitation by introducing an automatic adaptation of the interpolation space (mesh) during training, driven by a posteriori error estimation. We propose a NN error estimator based on the strong PDE residual (which is straightforward to implement and can be accurately computed via Gaussian quadratures) and a standard FE Kelly indicator. These advances allow for a more comprehensive exploitation of the non-linear nature of NNs, thus enhancing the capability of the method to tackle problems with localised features. The minimisation of the loss function relies on a train, estimate, mark, adapt procedure, which is repeated till a certain level of accuracy is reached. We leverage hierarchically adapted non-conforming octree-based meshes and adaptive FE spaces with hanging node constraints to keep the conformity of the interpolation.
@@
We conduct a comprehensive set of numerical experiments, focusing on forward problems featuring steep gradients in 2D and singularities in both 2D and 3D. We verify that expressive enough NNs can replicate the rate of convergence of the $h$-adaptive FEM solutions. In addressing the sharp-gradient problem, the generalisation capabilities of the trained NNs yield remarkable results. The non-interpolated NN solutions can be orders of magnitude more accurate than their interpolation counterparts on the same mesh. This observation is consistent with the findings in our previous work [Badia2024] for smooth solutions. Our findings reveal that the performance of the NN error indicator is on par with, if not surpassing, the traditional Kelly FE error estimator in various situations. Moreover, we confirm that a well-posed residual norm in the loss function has a tremendously positive impact on the convergence of the $h$-adaptive training process.
@@ 2.2. Finite element method
Problem (5) is well-posed under the assumption that the trial and test FE spaces $U_{h}$, $V_{h}$ have the same dimension and satisfy the following inf-sup condition [Griffiths1978]: there exists a constant $\beta>0$ independent of the mesh size $h$ such that
@@ 2.4. Finite element interpolated neural networks
We refer to the solution of (11) as an $\epsilon$-quasiminimiser of the FEINN problem. The concept of quasiminimisers in the frame of PINNs has already been previously discussed [Shin2023, Brevis2022]. This is of relative practical relevance since the problem is solved iteratively up to a certain level of accuracy and one can produce a sequence of $\epsilon$-quasiminimisers that converges to one of the solutions of the FEINN problem (10).
@@
It has been proven in [Petersen2020] that the space of functions that can be generated by a NN is not closed in $L^{p}$-norms for $0<p<\infty$ for most activation functions. For the $L^{\infty}$-norm, and (parameterised) ReLU activation functions, the NN space is proven to be closed in some cases. ReLU activation functions are not smooth enough for PINNs or Deep Galerkin methods [Magueresse2024]. However, in FEINNs, $\mathcal{N}_{h}$ only depends on the pointwise values of the neural network on the nodes of $U_{h}$ and ReLU activation functions are suitable, since the method computes spatial derivatives of FE functions. Thus, using the results in [Petersen2020, Sec. 3.2], one can check that $\mathcal{N}_{h}$ is closed for ReLU activation functions and the minimum in (10) is attained. This is not the case for other activation functions (see Ex. 3.11).
As discussed in [Badia2024], we can define a discrete Riesz projector $\mathcal{B}_{h}^{-1}:V_{h}^{\prime}\to V_{h}$ such that
@@
Unlike loss functions that rely on discrete norms of the residual vector, this loss function, preconditioned by $\mathcal{B}_{h}$, is well-defined in the limit $h\downarrow 0$. In practice, $\mathcal{B}_{h}^{-1}$ can be replaced by any spectrally equivalent approximation to reduce computational costs; e.g. one geometric multigrid (GMG) cycle has been proposed in [Badia2024].
@@
where $\left\lVert\cdot\right\rVert_{\chi}$ is an algebraic norm of a vector. The algebraic norm is ill-posed at the continuous level and, as a consequence, the convergence of the minimisation process is expected to deteriorate as $h\downarrow 0$ [Mardal2010].
@@ 2.6. Gradient-conforming discretisation
We consider an exotic Petrov-Galerkin FE method. The well-posedness of the discretisation is guaranteed by the discrete inf-sup condition (6) and the same dimension for trial and test spaces (see [Griffiths1978]). We use a grad-conforming nodal Lagrangian FE space $U_{h}$ of order $k_{U}$ built upon a (possibly non-conforming) partition $\mathcal{T}_{h}$ of $\Omega$. The FE space $U_{h}$ is built upon the partition $\mathcal{T}_{h}$. We use a linearised test FE space $V_{h}$ ($k_{V}=1$), built upon the partition $\mathcal{T}_{h}$ obtained after $k_{U}$ levels of uniform refinement applied to $\mathcal{T}_{h}$. It is easy to check that the dimension of ${U}_{h}$ and $V_{h}$ are identical. The reason behind this choice of the test space is the dramatic improvement in the convergence of the optimisation process compared to a Galerkin discretisation, as observed, e.g., in [Berrone2022, Badia2024].
@@ 2.7. Error indicators
The first error indicator is the classic Kelly error estimator [Kelly1983, Ainsworth1997], a well-established indicator in $h$-adaptive FEM, particularly for the Poisson equation. This error indicator estimates the error per element by integrating the gradient jump of the NN interpolation along the faces of each element. Specifically, for the element $K$, the Kelly error indicator is defined as
@@
The superscript $n$ in $\zeta_{K}^{n}$ stands for “network”. We note that $\zeta_{K}^{n}$ is similar to the mean PDE residual $\varepsilon_{r}$ defined in [Lu2021]. The main difference is that $\zeta_{K}^{n}$ is evaluated with a high-degree Gauss quadrature, while $\varepsilon_{r}$ is computed by Monte Carlo integration. At each mesh adaptation step, the strategy in [Lu2021] adds more collocation points only on the region with the highest $\varepsilon_{r}$. Instead, our approach can refine multiple elements with the largest errors and coarsen those with the smallest errors simultaneously.
@@ 3. Numerical analysis
In this section, we extend the analysis in [Badia2024] in multiple ways. In particular, in Sec. 3.1, we bound the error of the interpolated NN under the general scenario in which the NN cannot exactly emulate the FE space $U_{h}$, i.e., $\mathcal{N}_{h}\subsetneq U_{h}$ (or, equivalently, the interpolation operator $\pi_{h}:\mathcal{N}\mapsto U_{h}$ is not surjective). Remarkably, we prove that the error of the interpolated NN is bounded by the interpolation error among all possible vector spaces that result from the interpolation of the last-layer functions in the NN architecture. In Sec. 3.2, we perform error analysis under a weakened assumption of exact emulation which we refer to as quasi-emulation. This analysis is relevant as exact emulation might not be feasible for some activation functions, regardless of the expressivity of the NN. Finally, in Sec. 3.3, we also discuss the well-posedness of the FEINN problem in a quotient space and discuss the accuracy of the non-interpolated NN by considering the maximum distance between elements of an equivalence class.
@@ 3.1. Error analysis
First, we note that the loss function is differentiable (see [Badia2024, Prop. 3.1]). To obtain (17)-(18), we compute the variational derivative of (16) to variations of $u_{h}$ and recall the definition of the discrete Riesz projector. Under the assumption of the discrete inf-sup condition (6), we can prove the well-posedness of the problem using the Babuska-Brezzi theory. Namely, using (6), we can readily find $k_{h}\in V_{h}$ such that $\|k_{h}\|_{U}=\|u_{h}\|_{U}$ and $a(u_{h},k_{h})\geq\beta\|u_{h}\|_{U}^{2}$. We can take $v_{h}=r_{h}+\alpha k_{h}$ for an arbitrary $\alpha>0$ and $w_{h}=-u_{h}$, to get